Orange Inc. is considering two mutually exclusive projects (i.e., can choose either one but not both),Alpha in Country A, and Beta in Country B. Project Alpha requests an initial investment of $300,000 and has projected annual cash flows of $20,000, $50,000, $50,000 and $350,000 over 4 years. Project Beta requests an initial investment of $88,000 and has projected cash flows of $12,000 for the first year, and then the cash flows are projected to grow at a constant rate of 5 percent per year forever. Based on the project characteristics, the company requires an 15 percent return for Project Alpha but requires an 17 percent for Project Beta. You require a 15 percent return on your investment and a payback period of 3 years.

If you apply the NPV criterion, which investment will you choose? Why? (5 marks)
Based on your answer in (a) and (b), which project will you finally choose? Why? (3 marks)
1
Expert's answer
2019-12-10T08:43:21-0500

"NPV(A) = -300,000 + 20,000\/1.15 + 50,000\/1.15^2 + 50,000\/1.15^3 + 350,000\/1.15^4 = -11,812.06."

"NPV(B) = -88,000 + 12,000\/1.17 + 12,000\/(0.17 - 0.05) = 22,256.41."

So, you should choose project B.